the radius of the planet Mars is measured to be 3350 km , with a measurement error of plus or minus 100 km. What is the percent error when calculating the volume of the planet? Round your answer to 2 decimal places


1
Expert's answer
2021-12-20T03:26:39-0500

"Radius = 3350±100\\
\pi= 3350\\
Δ \pi= 100\\
V= \frac{4}{3} \pi r^2\\
\frac{Δ V}{V}*100=3(\frac{Δr}{r} *100)\\
\% error \space in \space V= \frac{Experimental \space value-Accepted \space value}{Experimental \space value}\\
= 3*\frac{100}{3350}*100\\
= 8.95 \%"